OpenStudy (anonymous):

Patti’s monthly pay is $3461.04. Each month she uses 2/3 of her pay for her household expenses and 1/8 for entertainment. She deposits the rest into her savings account. How much does Patti deposit into her savings account each month? $

OpenStudy (jdoe0001):

she'd end up depositing 3,461.04 - (2/3 of 3,461.04 + 1/8 of 3,461.04) or \(\bf 3,461.04-\cfrac{2}{3}3,461.04-\cfrac{1}{8}3,461.04 \\ \quad \\ 3,461.04-\left(\cfrac{2}{3}3,461.04+\cfrac{1}{8}3,461.04\right) \\ \quad \\ 3,461.04-\left(\cfrac{2\cdot 3,461.04}{3}+\cfrac{1\cdot 3,461.04}{8}\right) \\ \quad \\ 3,461.04-\left(\cfrac{6,922.08}{3}+\cfrac{ 3,461.04}{8}\right)\)